OpenStudy (anonymous):

Which of the following pairs of coordinates are on a line with an undefined slope? A. (-4, -3), (3, -3) B. (-2, -2), (4, 4) C. (2, 1), (5, -2) D. (3, -2), (3, 5)

Directrix (directrix):

D. (3, -2), (3, 5)

OpenStudy (anonymous):

D. (3, -2), (3, 5) m = (5 + 2) / (3 -3)= 7/0 = infinitive